Loli:

Which expression is equivalent to p^24 q^4/ p^8 q^16, if p ≠ 0 and q ≠ 0? A~ p^3 q^4 B~ p^3/q^4 C~ p^16 q^12 D~ p^16/q^12

Vocaloid:

consider p and q separately rule for exponents: x^m/x^n = x^(m-n)

Vocaloid:

p^(24)/p^8 = ?

Loli:

p^16

Vocaloid:

awesome what about q^4/q^16?

Loli:

1/q^12

Vocaloid:

awesome, let's put those together

Vocaloid:

p^16/q^12 which matches answer choice D

Loli:

Thank you! Again.